An oval track is made by erecting semicircles on each end of a 58m by 116m rectangle. Find the length of the track and area enclosed by the track. Use 3.14 for pie.
The length of the track is?
The area enclosed by the track is?

Answer by checkley77(12844)

IF the semicircle is on the 58m end then:
L=2*116+3.14*58
L=232+3.14*58
L=232+182.12=414.12 m. ans. for the length.
A=58*114+3.14(58/2)^2
A=6,612+3.14(29)^2
A=6,612+3.14*841
A=6,612+2,640.74
A=9,252.74 m^2. ans.
--------------------------
IF the semicircircle is on the 116m end then just replace the 58 by 116.